if sin(2a)=4sin2b,show that 5tan(a-b)=3tan(a+b)

Grade:12

1 Answers

Sunil Raikwar
9 years ago
sin2a/sin4b=4
sin2a+sin2b/sin2a-sin2b=4+1/4-1[by componendo & dividendo]
2sin(a+b)cos(a-b)/2cos(a+b)sin(a-b)=5/3
tan(a+b)/tan(a-b)=5/3
5tan(a-b)=3tan(a+b)
Thanks & Regards
Sunil Raikwar
askIITians Faculty